# A Simple yet Powerful Instance-Aware Prompting Framework for Training-free Camouflaged Object Segmentation

### v2

@@ Introduction
Ablation Study
Effectiveness of the Modules
We perform ablation studies to examine the effectiveness of each key component in the IAPF framework: Text Prompt Generator (TPG), Grounding DINO (G), Single-Foreground Multi-Background Prompting (SFMBP), and Self-consistency Instance Mask Voting (SIMV). As shown in Table 3, removing TPG (setting 1), G and SFMBP (settings 2.1-2.3), or SIMV (setting 3) clearly degrades performance, verifying the necessity of each module. Moreover, comparisons between settings (2.1 vs. 2.3) and (2.2 vs. 4) further demonstrate that introducing Grounding DINO not only grants the IAPF the critical instance-awareness capability but also significantly improves overall segmentation accuracy by allowing the model to localize more camouflaged instances. The complete method (setting 4) consistently outperforms all variants on both NC4K and COD10K-TEST datasets, confirming that each proposed component contributes distinctly and effectively to the final COS performance.
Effectiveness of Repetition Number
Performance Comparison of Different Camouflaged Instance Counts
The proposed IAPF maintains consistently high accuracy as instance density grows. The larger percentage swing in $M$ for all methods is expected because the absolute MAE values are very small; nonetheless, the absolute MAE of the IAPF remains the lowest. These results confirm that the instance-aware prompting pipeline––particularly the Grounding DINO–based box prompts and SFMBP point sampling––provides strong resilience to cluttered scenes with many camouflaged objects.
Conclusion
We introduced the IAPF, the first training-free framework that converts a single task-generic prompt into instance-level guidance for SAM, enabling zero-shot training-free segmentation of multiple camouflaged objects. By coupling Grounding DINO-based instance-level box prompts and the SFMBP point-sampling strategy, the IAPF delivers fine-grained COS without any additional training or annotations. Extensive experiments on standard COS benchmarks show that our method consistently outperforms prior unsupervised, weakly-supervised, and training-free COS approaches.

### v3

@@ 4.4 Downstream Applications
Methods NC4K COD10K-TEST
GenSAM .827 .774 .054 .880 .811 .721 .041 .867
ProMaC .826 .779 .053 .890 .818 .738 .038 .887
RDVP-MSD .823 .764 .056 .873 .825 .743 .038 .877
Ours .862 .828 .043 .916 .856 .799 .033 .917
Table 6: The fair comparison between the proposed method and other SOTA methods.
Methods ACOD-12K PlantCamo
GenSAM .478 .202 .193 .517 .670 .540 .114 .743
ProMaC .457 .175 .225 .499 .651 .521 .124 .757
RDVP-MSD .453 .147 .202 .505 .704 .582 .100 .761
ours .603 .406 .103 .660 .758 .664 .086 .829
Table 7: Downstream quantitative results on ACOD-12K and PlantCamo.
Figure 4: Downstream qualitative results on ACOD-12K and PlantCamo.
5 Conclusion
We presented IAPF, the first training-free COS framework that explicitly converts a single task-generic prompt into instance-level visual prompts—multiple boxes plus per-box foreground/multi-background points—for SAM. A frozen MLLM yields image-specific category tags; a detector-agnostic enumerator produces instance boxes; SFMBP derives contrastive per-box points; and SIMV selects the most self-consistent instance solution. These upgrades prompt granularity from semantic to instance-level and enable fine-grained, multi-instance COS without any task-specific training or annotations. Extensive results on three COS benchmarks, two CIS benchmarks, and two downstream datasets demonstrate state-of-the-art performance among training-free methods and strong cross-domain generalization, establishing instance-aware prompting as a simple yet powerful paradigm for promptable segmentation.